OpenStudy (anonymous):

A computer company gives a discount of 10 % on a new piece of software. If it makes a profit of 7%, what percent does the marked price exceed the cost price?

OpenStudy (anonymous):

let the cost price be x and marked y. y*90/100 =x*107/100 y/x=107/90 (y-x) 170 -----*100% = ----- = 18.89% x 9